# Pierre Dantas Dept. of Computer Science

### v1

@@ 5 Methodology
Our methodology establishes a comprehensive framework for validating the convergence theory through systematic simulation and analysis. We designed this approach to test the theoretical bounds rigorously; it simultaneously characterizes the behavior of LLM-verifier systems across their full operational range.
@@ 5.1 Simulation Framework Design
We developed a specialized simulation framework that implements the five-state Markov Chain model described in Section 4. The simulator implements probabilistic transitions between states Unverified, CandidateGenerated1, CandidateGenerated2, CandidateGenerated3, and Verified, with $\delta$ governing the success probability.
We summarize the complete experimental workflow in Figure 5. The process begins with the simulation core, which translates the five-state absorbing Markov Chain model (Section 4) into a highly efficient VMC simulation (Section 5.2.2).
We examine the success probability parameter $\delta$ across its entire range. We categorize the resulting operational data into three key regions: marginal ($\delta\in[0.1,0.3)$), practical ($\delta\in[0.3,0.6]$), and high-performance ($\delta\in(0.6,0.9]$). After generating the large-scale dataset, we proceed to the data collection phase. In this stage, we record both the total convergence time ($T$) and the sequence traces. The final stage involves the experimental validation framework, which uses a three-pronged analytical approach – theoretical, sensitivity, and statistical – to rigorously test the derived bound $\mathbb{E}[n]\leq 4/\delta$ and to provide empirically validated engineering insights.
Figure 5: Overview of the LLM-Verifier Convergence validation methodology. The process links the theoretical five-state absorbing Markov Chain model to the experimental framework. The simulation core uses a VMC approach with the Geometric distribution ($\text{Geom}(\delta)$) to efficiently generate $N=10,000$ trials across the three defined operational regions (marginal, practical, high-performance). After data collection (generating convergence time $T$ and computational metrics), the results are channeled into the experimental validation framework, which employs three complementary analytical lenses to confirm theoretical alignment, assess framework robustness, and ensure statistical reliability (see Table 7)
@@
      The marginal region $(\delta\in[0.1,0.3))$ addresses small success probabilities where convergence is theoretically ensured by exponential tail bounds, $\mathbb{P}(n>k)\leq e^{-c\delta k}$ [Boucheron2013, Papaspiliopoulos2020], but tests practical computational limits.
@@
      Finally, the high-performance region $(\delta\in(0.6,0.9])$ represents near-expert behavior, corresponding to the high-probability convergence regime in Monte Carlo and reliability theory [Robert2010, Wasserstein1997].
@@ 5.2 Simulation Execution
    Geometric Distribution: The Geometric distribution, $\text{Geom}(\delta)$, models the time spent in each of the four transient states ($s_{1}$ through $s_{4}$). Its expected value, $1/\delta$, directly establishes the theoretical basis for the primary bound we are validating: $\mathbb{E}[n]=\sum_{j=1}^{4}\mathbb{E}[M_{j}]=4/\delta$.
@@
    • Sequence of state transitions: Enables analysis of the convergence path, identifying patterns and bottlenecks to verify the modeled Markovian dynamics.
    • Convergence success/failure status: Provides binary outcomes to empirically verify the Theorem’s guarantee of almost sure convergence across all trials.
@@ 5.2.3 Data Collection and Filtering
  • Sequence Trace: Saving the complete list $\mathbf{M}_{i,1},\mathbf{M}_{i,2},\mathbf{M}_{i,3},\mathbf{M}_{i,4}]$ for each trial enables us to identify specific patterns where progress slows. For example, we can look at trials that went through the first three steps quickly but got stuck at the last step of checking.
@@ 6.2.1 Internal Validity
  • Parameter Stationarity: This means we assume the parameter $\delta$ stays the same during improvement. In reality, LLMs can have varying performance over time.
  • State Space Limitation: Our model has five states, which may not capture all the details needed for complex tasks that require checking.
@@ 6.2.2 External Validity
  • Simulation vs. Real Systems: Experiments use modeled behavior (sum of Geometric distributions); validation with real LLM and application programming interfaces (APIs) is future work.
  • Domain Generalization: Results based on synthetic Markov Chains; domain-specific $\delta$ calibration needed for different verification tasks.
@@ 6.2.3 Construct Validity
  • Success Definition: Binary success/failure may not capture partial progress in complex refinements.
@@ 7 Discussion
This section interprets the empirical results presented in Section 6.1 against the theoretical model and core objectives established in Section 6, providing actionable insights for the design of LLM-verifier systems.
@@ 7.1 Validation of Theoretical Bound and Alignment (RQ2)
The empirical validation decisively confirms the theoretical bound $\mathbb{E}[n]\leq{4}/{\delta}$ as a predictable ceiling for convergence time. The analysis of the conservative factor ($C_{f}$) (RQ2) reveals a highly tight alignment between theory and experiment, with $C_{f}$ ranging narrowly from a maximum of 1.0043 to a minimum of 0.9942 (Table 8, p. 8). This closeness confirms that the theoretical bound perfectly predicts the empirical mean convergence time. This finding provides a critical engineering guarantee, elevating the determination of resource allocation from empirical guesswork to a practice governed by a robust, verifiable mathematical guarantee.
@@ 7.2 Phase Transition and System Predictability (RQ3 & RQ5)
Looking at the variance ($\sigma^{2}$) helps us confirm the set areas where operations can happen (RQ3) and shows how the entire range of values changes over time (RQ5). The way the standard deviation decreases quickly supports the idea that there is an important change in how predictable the system is.
@@
  • This validates the defined $\delta$ thresholds: systems entering the practical region mitigate the variance risk inherent to low-$\delta$ configurations, which is essential for reliable service-level adherence.
This data shows that when we improve the abilities of LLMs, we see big improvements in how well they can predict outcomes. We especially want to focus on the area where performance is high (when the value is greater than 0.6) because that area is important for creating systems that are reliable and have little variation.
@@ 7.3 Alignment with Theoretical Premises (RQ1 & RQ4)
The results of the experiments align perfectly with the study’s basic ideas, confirming the main arguments about how things come together.
  • Almost-Sure Convergence (RQ1): The 100% success rate shows strong evidence that the system called the LLM-verifier works well and can fix itself. This is true as long as a certain condition (called $\delta$) is met. The way it operates is based on specific patterns of behavior, similar to how a Markov Chain works.
  • Statistical Distribution Fit (RQ4): The structure of the empirical data strongly confirms that the convergence time follows a negative binomial distribution. The steady decrease seen in the tail probability analysis (Figure 8) supports the main mathematical model used for the state-transition process. This result effectively connects the unpredictable behavior of probabilistic LLMs with the set rules needed in a more certain way.
@@ 7.4 Computational Feasibility and Framework Validity (RQ6)
The simulation framework’s performance validates its utility as a high-speed analytical tool for LLM assessment (RQ6). A throughput exceeding 858,000 trials/second, achieved with minimal memory overhead, demonstrates that the analytic approach is highly scalable and computationally feasible. This performance validates the framework’s ability to support continuous monitoring and dynamic testing of LLM capability in production environments.
@@ 8 Conclusion and Future Work
The experimental validation decisively confirms the theoretical premises and core objectives of this work. The detailed simulation of the VMC and LLM-verifier system showed an intense match with the proposed mathematical rules. This proves that the convergence theory provides a solid basis for combining Formal Methods and AI.
The results clearly showed that our initial idea of almost-sure convergence was correct (RQ1), with a 100% success rate across all 90,000 tests. Importantly, we also confirmed the expected limit of iterations, which is $\mathbb{E}[n]\leq 4/\delta$, with great accuracy (RQ2). The average results closely matched our predictions, and we found that the conservative factor ($C_{f}$) remained consistently within the range $[0.9942,1.0043]$ across all tests. This consistency proves that our theoretical model is accurate in different performance conditions.
The statistical analysis showed precise numbers that helped define three different operational regions (RQ3). The variation in data ($\sigma^{2}$) changed significantly, dropping from about 18.84 in the marginal region ($\delta=0.1$) to around 0.71 in the high-performance region ($\delta=0.9$). This significant change supports the idea of separating these regions.
Looking at the complete statistical distribution (RQ5) and the tail probabilities (RQ4) confirmed that the convergence process matches the expected patterns, indicating that the worst-case scenarios are reliable based on the statistics. Finally, the analytical framework proved highly scalable and applicable (RQ6), capable of running more than 858,000 trials per second.
In summary, the experimental results provide direct, strong, and affirmative confirmation that the convergence theory provides the mathematical foundation for the integration and reliable deployment of the formal LLM method.
@@ 8.1 Future Work
The theories and validation methods created in this research point to many exciting opportunities for future studies. These mainly focus on turning simulation results into real-life uses and making the mathematical model more detailed and complex.
A critical future direction is the full integration of the framework with LLM, which involves extending the current simulation environment to interface directly with the LLM’s APIs. The main goal is to measure how well a model performs in different verification tasks. We want to take the theoretical number $\delta$ and turn it into something we can actually measure about how capable the model is when it is used. To do this, we need to adjust it based on the specific area we are working in. This means we will closely study how the success rate, represented by $\delta$, changes in different verification areas and as tasks become more or less complex.
In terms of mathematical rigor, future work should focus on advanced modeling, such as developing continuous-state and time-varying capability models to capture more nuanced and complex refinement dynamics. Such models would move beyond the discrete, constant probability $\delta$ and account for learning or fatigue effects over the course of the iterative process. Finally, the insights gained from empirical $\delta$ measurement will enable the creation of adaptive systems. These systems can dynamically adjust their verification strategy, resource allocation, or prompt structure based on a real-time estimate of the LLM’s current success probability ($\delta$).
@@ 8.2 Concluding Remarks
This work has created a strong basis for understanding how LLMs can work together with Formal Methods. By achieving our research goals, we have provided solid mathematical support for a process that used to rely on guesses. This change turns a trial-and-error approach into a more reliable and precise way of engineering. The key findings directly reaffirm the abstract and objectives:
  • The empirical 100% success rate validated the almost-sure convergence proof.
  • The theoretical bound $\mathbb{E}[n]\leq{4}/{\delta}$ was confirmed to exhibit tight alignment ($C_{f}\approx 1.0$), demonstrating the accuracy of the prediction (RQ2).
  • The empirical data successfully defined three distinct operational regions, based on the exponential drop in variance from $\sigma\approx 18.84$ to $\sigma\approx 0.71$ (RQ3/RQ5).
@@
This framework establishes a new paradigm for human-AI collaboration in critical reasoning tasks, seamlessly integrating the complementary strengths of learning and logic to solve the most challenging problems in computer science and beyond.

### v2

@@ 5 Methodology
Our methodology establishes a comprehensive framework for validating the convergence theory through systematic simulation and analysis. We designed this approach to test the theoretical bounds rigorously; it simultaneously characterizes the behavior of sequential LLM-verifier systems across their full operational range.
@@ 5.1 Simulation Framework Design
We developed a specialized simulation framework that implements the sequential Markov Chain model described in Section 4. The simulator implements probabilistic transitions through the four distinct pipeline stages: CodeGen ($s_{1}$), Compilation ($s_{2}$), InvariantSynth ($s_{3}$), and SMTSolving ($s_{4}$), leading ultimately to Verified ($s_{5}$). In this model, $\delta$ governs the success probability of advancing from stage $s_{i}$ to $s_{i+1}$.
We examine the success probability parameter $\delta$ across its entire range. We categorize the resulting operational data into three key regions: marginal ($0.1\leq\delta<0.3$), practical ($0.3\leq\delta\leq 0.6$), and high-performance ($0.6<\delta\leq 0.9$). After generating the large-scale dataset, we proceed to the data collection phase. In this stage, we record both the total convergence time ($T$) and the sequence traces. The final stage involves the experimental validation framework, which uses a three-pronged analytical approach – theoretical, sensitivity, and statistical – to rigorously test the derived bound $\mathbb{E}[n]\leq 4/\delta$ and to provide empirically validated engineering insights.
We summarize the complete experimental workflow in Figure 5. The process begins with the simulation core, which translates the sequential pipeline model (Section 4) into a highly efficient VMC simulation (Section 5.2.2).
Figure 5: Overview of the LLM-Verifier Convergence validation methodology. The process links the theoretical sequential Markov Chain model to the experimental framework. The simulation core uses a VMC approach with the Geometric distribution ($\text{Geom}(\delta)$) to efficiently generate $N=10,000$ trials across the three defined operational regions. After data collection (generating convergence time $T$ and computational metrics), the results are channeled into the experimental validation framework, which employs three complementary analytical lenses to confirm theoretical alignment, assess framework robustness, and ensure statistical reliability (see Table 7)
@@
      The marginal region $(0.1\leq\delta<0.3)$ addresses small success probabilities where convergence is theoretically ensured by exponential tail bounds, $\mathbb{P}(n>k)\leq e^{-c\delta k}$ [Boucheron2013, Papaspiliopoulos2020], but tests practical computational limits.
@@
      Finally, the high-performance region $(0.6<\delta\leq 0.9)$ represents near-expert behavior, corresponding to the high-probability convergence regime in Monte Carlo and reliability theory [Robert2010, Wasserstein1997].
@@ 5.2 Simulation Execution
    Geometric Distribution: The Geometric distribution, $\text{Geom}(\delta)$, models the residence time spent in each of the four pipeline stages ($s_{1}$ through $s_{4}$). Its expected value, $1/\delta$, directly establishes the theoretical basis for the primary bound we are validating. Since the process is sequential, the total expected time is the sum of the means: $\mathbb{E}[n]=\sum_{j=1}^{4}\mathbb{E}[M_{j}]=4/\delta$.
@@
    • Sequence of state transitions: Enables analysis of the convergence path, identifying bottlenecks in specific pipeline stages (e.g., verifying if Invariant Synthesis $s_{3}$ consumes disproportionate resources).
    • Convergence success/failure: Provides binary outcomes to empirically verify the Theorem’s guarantee of almost sure convergence across all trials.
@@ 5.2.3 Data Collection and Filtering
  • Sequence Trace: Saving the complete list $\mathbf{M}_{i,1},\mathbf{M}_{i,2},\mathbf{M}_{i,3},\mathbf{M}_{i,4}]$ for each trial enables us to identify specific patterns where progress slows. For example, we can identify trials that clear the initial CodeGen and Compilation stages quickly but stall during the final SMT Solving stage.
@@ 6.2.1 Internal Validity
Internal validity assesses whether the study accurately demonstrates a causal relationship between the variables, ensuring that the observed results are attributable to the proposed theoretical framework rather than artifacts of the experimental design.
  1. 1.
@@
  2. 2.
    Parameter Stationarity: We assume the parameter $\delta$ remains constant throughout the refinement process to derive a closed-form bound. In reality, LLMs may exhibit non-stationary behavior (e.g., performance drift due to context window saturation). While this abstraction simplifies the model, it represents the “average effective capability” of the system. Future work (discussed in Section 7.5) will address dynamic calibration strategies to mitigate this drift in real-world deployments.
  3. 3.
    State Space Limitation: Our model has five states, which captures the core engineering pipeline ($s_{1}\to s_{5}$) but may not capture granular sub-states within complex tasks (e.g., partial invariant synthesis).
  4. 4.
@@ 6.2.2 External Validity
External validity evaluates the generalizability of our findings beyond the controlled simulation environment. This section assesses the extent to which the theoretical bounds and operational regions derived from our model apply to real-world LLM deployments, diverse verification domains, and varying engineering scales.
  • Simulation vs. Real Systems: Experiments use modeled behavior (sum of independent Geometric distributions); validation with real LLM and application programming interfaces (APIs) is future work.
  • Domain Generalization: Results based on synthetic Markov Chains; domain-specific $\delta$ calibration is needed for different verification tasks (e.g., SMT solving vs. proof synthesis).
@@ 6.2.3 Construct Validity
Construct validity examines whether the operational measures used in the study essentially represent the theoretical constructs they are intended to measure. We analyze potential disconnects between our mathematical metrics and the real-world engineering goals of LLM verification.
@@
  • Success Definition: Binary success/failure provides a clear convergence signal but may not capture partial progress in complex refinements.
@@ 7 Discussion
This section interprets the empirical results presented in Section 6.1 against the theoretical model and core objectives established in Section 6, providing actionable insights for the design of sequential LLM-verifier systems.
@@ 7.1 Validation of Theoretical Bound and Alignment (RQ2)
The empirical validation decisively confirms the theoretical bound $\mathbb{E}[n]\leq{4}/{\delta}$ as a predictable ceiling for convergence time. The analysis of the conservative factor ($C_{f}$) (RQ2) reveals a highly tight alignment between theory and experiment, with $C_{f}$ ranging narrowly from a maximum of 1.0043 to a minimum of 0.9942 (Table 8, p. 8).
This closeness confirms that the total verification latency is indeed the sum of the latencies of the four pipeline stages ($s_{1}\to s_{4}$). This finding provides a critical engineering guarantee: it elevates the determination of resource allocation from empirical guesswork to a practice governed by the robust additive logic of the sequential pipeline, where total cost is predictably $\sum\mathbb{E}[M_{j}]=4/\delta$.
@@ 7.2 Phase Transition and System Predictability (RQ3 & RQ5)
Looking at the variance ($\sigma^{2}$) helps us confirm the operational regions (RQ3) and shows how the system stability evolves (RQ5). The rapid collapse of the standard deviation supports the existence of a critical phase transition in system predictability.
@@
  • This validates the defined $\delta$ thresholds: systems entering the practical region mitigate the variance risk inherent to low-$\delta$ configurations. For engineers, this implies that optimization efforts should focus on pushing $\delta$ past the $0.3$ threshold to secure reliable service-level adherence.
This data demonstrates that improving LLM capability does not just reduce average time; it fundamentally alters the stability of the pipeline. We especially identify the High-Performance region ($\delta>0.6$) as the target for safety-critical systems requiring minimal variance.
@@ 7.3 Alignment with Theoretical Premises (RQ1 & RQ4)
The experimental results align perfectly with the study’s fundamental premises, mathematically validating the structure of the pipeline.
  • Almost-Sure Convergence (RQ1): The 100% success rate provides strong evidence that the sequential LLM-verifier is statistically guaranteed to self-correct and terminate. This holds true provided that the probability of success for any single stage ($\delta$) remains non-zero.
  • Statistical Distribution Fit (RQ4): The structure of the empirical data strongly confirms that the convergence time follows a Negative Binomial distribution with parameter $r=4$. Since a Negative Binomial($r,p$) distribution describes the sum of $r$ independent Geometric variables, this empirical fit mathematically validates our modeling assumption that the process consists of exactly four independent, sequential refinement stages ($s_{1},s_{2},s_{3},s_{4}$).
@@ 7.4 Computational Feasibility and Framework Validity (RQ6)
The simulation framework’s performance validates its utility as a high-speed analytical tool for LLM assessment (RQ6). A throughput exceeding 858,000 trials/second, achieved with minimal memory overhead, demonstrates that the analytic approach is highly scalable. This performance validates the framework’s ability to support continuous monitoring and dynamic testing of LLM capability in production environments.
7.5 Handling Non-Stationary Environments (Dynamic Calibration)
A key threat to validity identified in Section 6.2 is parameter stationarity – the risk that $\delta$ may drift over time due to LLM context saturation or fatigue. While our Theorem assumes a fixed $\delta$, our framework offers a practical mitigation strategy for real-world deployment: dynamic calibration.
Because the system behavior is well-characterized by the operational regions (Table 1, p. 1), engineers can implement a runtime monitor that calculates a sliding-window average of the success rate ($\hat{\delta}$).
  • If $\hat{\delta}$ drops into the marginal region ($\hat{\delta}<0.3$), the system can automatically trigger a context reset or increase the sampling temperature to restore the error-reduction probability.
  • This approach effectively treats a non-stationary process as a sequence of piecewise-stationary segments, ensuring that the system dynamically adheres to the convergence guarantees provided by the Theorem.
@@ 8 Conclusion and Future Work
The experimental validation decisively confirms the theoretical premises and core objectives of this work. The detailed simulation of the VMC and LLM-verifier system demonstrated a robust alignment with the proposed sequential absorbing Markov Chain model. This proves that modeling verification not as a generic loop, but as a structured pipeline of engineering stages ($s_{1}\to s_{4}$), provides a solid mathematical basis for combining Formal Methods and AI.
The results clearly showed that our initial hypothesis of almost-sure convergence was correct (RQ1), with a 100% success rate across all 90,000 tests. Importantly, we confirmed that the cumulative expected latency is governed by the sum of the individual stage latencies, validating the bound $\mathbb{E}[n]\leq 4/\delta$ with high accuracy (RQ2). The average results closely matched our predictions, and the conservative factor ($C_{f}$) remained consistently within the range $[0.9942,1.0043]$. This consistency proves that the linear accumulation of time across the four pipeline stages is a valid predictor of total system cost.
The statistical analysis provided precise metrics to define three different operational regions (RQ3). The variance in convergence time ($\sigma^{2}$) collapsed significantly, dropping from approximately 18.84 in the marginal region ($\delta=0.1$) to around 0.71 in the high-performance region ($\delta=0.9$). This phase transition supports the architectural recommendation to treat these regions differently in deployment.
Looking at the complete statistical distribution (RQ5) and the tail probabilities (RQ4) confirmed that the convergence process follows the predicted geometric decay, indicating that worst-case outliers are mathematically bounded. Finally, the analytical framework proved highly scalable (RQ6), capable of running more than 858,000 trials per second.
In summary, the experimental results provide direct, strong, and affirmative confirmation that the convergence theory provides the mathematical foundation for the integration and reliable deployment of sequential LLM verification pipelines.
@@ 8.1 Future Work
The theories and validation methods created in this research point to distinct opportunities for future studies, particularly in addressing the complexities of real-world deployment.
  • Integration with Real LLM APIs: A critical next step is extending the simulation environment to interface directly with LLM APIs. We aim to measure empirical $\delta$ values for specific stages (e.g., measuring $\delta_{\text{code}}$ vs. $\delta_{\text{invariant}}$) to validate if the success rate remains uniform across different phases of the pipeline.
  • Handling Non-Stationary Environments: As noted in our validity analysis, real-world LLMs may exhibit parameter drift (non-stationarity) due to context window saturation or fatigue. Future work will focus on developing dynamic calibration strategies. These adaptive systems will estimate $\delta$ in real-time using a sliding window; if performance drops below the practical region threshold, the system could dynamically trigger context resets or strategy shifts.
  • Advanced Probabilistic Modeling: We aim to refine the mathematical model by exploring continuous-time Markov Chains to capture nuanced timing effects beyond discrete steps. This includes modeling correlations between stages (e.g., where a weak success in $s_{1}$ increases the failure probability in $s_{2}$).
@@ 8.2 Concluding Remarks
This work has created a strong basis for understanding how LLMs can work together with Formal Methods. By achieving our research goals, we have provided solid mathematical support for a process that used to rely on heuristics. This change turns a trial-and-error approach into a verifiable engineering discipline. The key findings directly reaffirm the objectives:
  • The empirical 100% success rate validated the almost-sure convergence proof for the sequential pipeline.
  • The theoretical bound $\mathbb{E}[n]\leq{4}/{\delta}$ was confirmed to exhibit tight alignment ($C_{f}\approx 1.0$), demonstrating that total latency is predictable based on the four-stage structure (RQ2).
  • The empirical data successfully defined three distinct operational regions, identifying a critical stability phase transition at $\delta\approx 0.3$ (RQ3/RQ5).
@@
This framework establishes a new paradigm for human-AI collaboration in critical reasoning tasks, seamlessly integrating the complementary strengths of probabilistic learning and sequential logic to solve challenging problems in computer science.